OpenStudy (anonymous):

Celina measured the height of a window frame as 5.55 feet, but the actual height was 6 feet. What is the percentage of error in Celina's measurement? 6.45% 7.50% 9.12% 13.33%

jimthompson5910 (jim_thompson5910):

https://www.mathsisfun.com/numbers/percentage-error.html \[\large \text{percent error} = \frac{| \text{approximate value}-\text{exact value}|}{|\text{exact value}|} \ \times 100\%\]

OpenStudy (anonymous):

5.55 - 6/6 x 100?

jimthompson5910 (jim_thompson5910):

I think you meant to say \[\Large \text{percent error} = \frac{| 5.55-6|}{|6|} \times 100\%\] right?

OpenStudy (anonymous):

i got -94.45...

jimthompson5910 (jim_thompson5910):

what is 5.55 - 6 equal to?

OpenStudy (anonymous):

-0.45

jimthompson5910 (jim_thompson5910):

|-0.45| = 0.45

jimthompson5910 (jim_thompson5910):

divide that by 6 then multiply by 100 to convert to a percentage

OpenStudy (anonymous):

is it b?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

i got 7.50

jimthompson5910 (jim_thompson5910):

me too

OpenStudy (anonymous):

ok thank you :)

jimthompson5910 (jim_thompson5910):

7.50%